OpenStudy (anonymous):

Solve and check you're solution: 2/5n=3/10n-3/5

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

2/5 n = 3/10 n - 3/5 3/5 = 3/10 n - 2/5 n 3/5 = (3-4)/10 n 3/5 = -1/10 n n = -30/5 n = -6

OpenStudy (anonymous):

when I checked the problem it came out to be -11/20=-1/5

OpenStudy (anonymous):

i'm pretty sure n =-6

OpenStudy (gw2011):

The answer is n=-6 Here is the check: (2/5)(-6)=(3/10)(-6)-3/5 -12/5=-9/5-3/5 -12/5=-12/5